OpenStudy (anonymous):

Determine the equation of a tangent to the curve defined by the equation F(x): -7x+5 X =8 The slope-intercept form for the equation of the tangent line to curve X=8 , and Y is ? If you don't mind explaining how you get your answer, so I can learn. I really appreciate it thanks!

OpenStudy (agentjamesbond007):

equation of the tangent line is the derivative using rules of derivation, \[cx ^{n}\] --> \[cnx ^{n-1}\] f'(x) = 7 plug your x-coordinate to get the y since 7 is a constant, it is 0. So you're solving for the equation at (8,0) Use point-slope y-y1=m(x-x1) y-0=7(x-8) your equation of the slope of the curve at (8,0) or x=8 is y=7x-56

OpenStudy (anonymous):

thank you I know where I messed up, i solved as if it was a regular line. Ha!

OpenStudy (agentjamesbond007):

You're Welcome!